A steel bar will expand when heated and contract when cooled. The relationship between the length of the bar, L, (mm) and the temperature, T, (•C) is given by
L = 5000 + 0.12(T- 20)
What is the the of the steel bar when the temperature is 45•C

Answer by stanbon(75887)

L = 5000 + 0.12(T- 20)
What is the length of the steel bar when the temperature is 45•C
===============
L = 5000+ 0.12(45-20)
---
L = 5000+0.12(25)
L = 5000+3
L = 5003 mm
======================
Cheers,
Stan
=====================